Formulate a quadratic equation then answer what is asked in the problem.
If John plans to put a Computer Sales and Services with printing station and applies for a loan of an amount P in pesos and r percent compounded annually to invest in his business, he believes that his business will grow to an amount A=P(1+r)² in two years. Supposed John wants his money amounting to Php200,000 to grow to Php228,980 in two years. At what rate must he invest his money?

Answer by Theo(13342)

the formula to use if f = p * (1 + r) ^ n
f is the future value
p is the present value
r is the interest rate per time period
n is the number of time period.

in your problem, the formula becomes:
228,980 = 200,000 * (1 + r) ^ 2

divide both sides of the equation by 200,000 and take the second root of both sides of the equation and subtract 1 from both sides of the equation, in that order, to get:

(228,980/200,000) ^ (1/2) - 1 = r
solve for r to get:
r = .07

he must invest 200,000 at 7% per year, compounded annually, to have 228,980 in 2 years.

200,000 * (1.07) ^ 2 = 228,980.

7% per year is your answer.